# Nature of spinons in the 1D spin chains

### v2

@@ III Adding a single spin to other 1D spin model ground states
Figure 4: Dispersion relation $\varepsilon(q)$ of the excitation calculated using the same procedure as discussed in the main text of the paper for: (a-g) the XXZ model (34) with different values of the anisotropy $\Delta$, (h) the Majumdar-Ghosh model (35) that has the valence-bond solid as the ground state. Results for chains of various length $L$ are shown by symbols (the connecting dashed lines are guides to the eye). The thick black solid lines, for comparison purposes, show the exact analytical expressions: (a-g) for the XXZ model originating in the Bethe Ansatz, and (h) for the Majumdar-Ghosh model, see text for more details.
Here we show that the procedure, discussed in the main text works well also for other 1D spin models. [We note in passing a (successful) application of the procedure to the Shastry-Sutherland model discussed in Talstra and Strong (1997)]. To this end, we consider two models. The first is the XXZ model defined by
$H^{\rm XXZ}_{L}=J\sum_{j=1}^{L}\left(\sigma^{x}_{j}\cdot\sigma^{x}_{j+1}+\sigma^{y}_{j}\cdot\sigma^{y}_{j+1}+\Delta\sigma^{z}_{j}\cdot\sigma^{z}_{j+1}\right),$ (34)
where we set $J=1$ and widely vary the anisotropy $\Delta$ (cf. Talstra and Strong (1997) for the XY model result). The second is the Majumdar-Ghosh model with its Hamiltonian
$H^{\rm MG}_{L}=J\sum_{j=1}^{L}\left(\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+1}+\frac{1}{2}\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+2}\right),$ (35)
@@
The corresponding dispersions $\varepsilon(q)$ obtained through the procedure proposed in the main text, are shown in Fig. 4. For the XXZ model with easy axis anisotropy we observe perfect agreement with the results obtained using the Bethe Ansatz Caux et al. (2008); Franchini (2017); Bohrdt et al. (2018); Rutkevich (2022). The agreement is quantitatively worse for the XXZ model with an easy plane anisotropy: In this case more than one spinon is excited when a single spin is inserted in the ground state, cf. the XY limit discussed in Talstra and Strong (1997). For the Majumdar-Ghosh model only small deviations are observed from the exact dispersion, see Ref. Shastry and Sutherland (1981); Lavarélo and Roux (2014).
@@
As a side note, in the Majumdar-Ghosh model the dispersion scales as ∝$\propto q^{2}$ in the low-energy limit Shastry and Sutherland (1981); Bohrdt et al. (2018), despite having a ground state with valence bond character and thus the norm $N(q)$ nearly vanishing in an extended region of $|q|>\pi/2$, just as in Fig. 2(c2) of the main text. The conundrum is resolved by noting that in the Majumdar-Ghosh Hamiltonian Shastry and Sutherland (1981) there exist next-nearest-neighbor spin flip terms which lead to a ∝$\propto\cos(2q)$ contribution to the dispersion relation.

### v3

@@ III Adding a single spin to other 1D spin model ground states
Figure S2: Dispersion relation $\varepsilon(q)$ of the excitation calculated using the same procedure as discussed in the main text of the paper for: (a-g) the XXZ model (SIII.24) with different values of the anisotropy $\Delta$, (h) the Majumdar-Ghosh model (SIII.25) that has the valence-bond solid as the ground state. Results for chains of various length $L$ are shown by symbols (the connecting dashed lines are guides to the eye). The thick black solid lines, for comparison purposes, show the exact analytical expressions: (a-g) for the XXZ model originating in the Bethe Ansatz, and (h) for the Majumdar-Ghosh model, see text for more details.
Here we show that the procedure, discussed in the main text works well also for other 1D spin models. [We note in passing a (successful) application of the procedure to the Shastry-Sutherland model discussed in Talstra and Strong (1997)]. To this end, we consider two models:
The first is the XXZ model defined by
$H^{\rm XXZ}_{L}=J\sum_{j=1}^{L}\left(\sigma^{x}_{j}\cdot\sigma^{x}_{j+1}+\sigma^{y}_{j}\cdot\sigma^{y}_{j+1}+\Delta\sigma^{z}_{j}\cdot\sigma^{z}_{j+1}\right),$ (SIII.24)
where we set $J=1$ and widely vary the anisotropy $\Delta$ (cf. Talstra and Strong (1997) for the XY model result). Note that the XXZ model with $\Delta=-1$ is equivalent to the FM Heisenberg model (the unitary transformation that relates these models shifts the wavevector of the magnon dispersion minimum).
The second is the Majumdar-Ghosh model with its Hamiltonian
$H^{\rm MG}_{L}=J\sum_{j=1}^{L}\left(\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+1}+\frac{1}{2}\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+2}\right),$ (SIII.25)
@@
The corresponding dispersions $\varepsilon(q)$ obtained through the procedure proposed in the main text, are shown in Fig. S2. For the XXZ model with easy axis anisotropy we observe perfect agreement with the results obtained using the Bethe Ansatz Caux et al. (2008); Franchini (2017); Bohrdt et al. (2018); Rutkevich (2022). The agreement is quantitatively worse for the XXZ model with an easy plane anisotropy: In this case more than one spinon is excited when a single spin is inserted in the ground state, cf. the XY limit discussed in Talstra and Strong (1997). For the Majumdar-Ghosh model only small deviations are observed from the exact dispersion, see Ref. Shastry and Sutherland (1981); Lavarélo and Roux (2014).
@@
As a side note, in the Majumdar-Ghosh model the dispersion scales as $\propto q^{2}$ in the low-energy limit Shastry and Sutherland (1981); Bohrdt et al. (2018), despite having a ground state with valence bond character and thus the norm $N(q)$ nearly vanishing in an extended region of $|q|>\pi/2$, just as in Fig. 2(c2) of the main text. The conundrum is resolved by noting that in the Majumdar-Ghosh Hamiltonian Shastry and Sutherland (1981) there exist next-nearest-neighbor spin flip terms which lead to a $\propto\cos(2q)$ contribution to the dispersion relation.
IV Pump-probe experimental techniques that can excite a single spinon
We postulate that the following ‘pump-probe’ style experiment should be able to measure a single spinon following the procedure outlined in the paper. While as of now it can be seen merely as a ‘Gedanken’ experiment, it is conceivable that it can be realised in the future. The main concept behind this experiment is related to removing a single spin from a magnetic lattice and enhancing the spin exchange across the empty site in such away that it is indistinguishable from the spin exchange between the intact bonds – in short we just ‘kick a spin and then glue the spins together’. Note that this procedure is equivalent to removing a single spin site and ‘contracting’ the lattice – yet it does not require changing the geometry of the lattice and hence not only it is valid in any dimension but also should not so easily excite phonons.
In detail, such an experiment can be designed in the following manner:
  1. 1.
    At time $\tau=0$ and momentum $k$ in the 1st Brillouin zone a photoemission experiment (PES) creates a hole in the valence shell (i.e. one spin is kicked out with a well-defined momentum from the valence shell). This pump and its action is illustrated in left panel of Fig. S3.
  2. 2.
    At time $\tau\ll 1/t$ and momentum $k$ in the 1st Brillouin zone an indirect resonant inelastic x-ray scattering (RIXS) experiment (e.g. $K$ edge) creates a core hole. This way simultaneously with creating a hole in the spin lattice, a core-hole potential is created at the same site (in momentum space), see left panel of Fig. S3.
  3. 3.
    Next, at time $\tau\in[1/t,\Gamma]$, where $\Gamma$ is the core-hole lifetime, the system can create a single spinon excitation, provided the core-hole potential $U_{c}$ has a particular value w.r.t $U$:
    For instance, if we assume that $U\sim 8t$ (a rather standard value for a Hubbard model that yields the typical spin exchange $J=0.5t$) and that $U_{c}\sim U/3$ then we obtain that the spin exchange $\tilde{J}$ between the spins across the empty site  [^1] Here we apply the standard superexchange formula, [cfZaanen and Oleś (1988) or Eq. (7) in Wohlfeld et al. (2010)] to the bond with three transition metal ions. reads
    $\tilde{J}=4\frac{t^{4}}{\Delta^{2}}\left(\frac{2}{2\Delta+U}+\frac{1}{U}\right)\Big{|}_{\Delta=-U_{c}}=4\frac{t^{4}}{U_{c}^{2}}\left(\frac{2}{U-2U_{c}}+\frac{1}{U}\right)\Big{|}_{U_{c}=U/3\ \&\ U=8t}\sim 0.5t,$ (SIV.26)
    and hence it is equal to the ‘canonical’ spin exchange $J=0.5t$ between all ‘other’ spins, see middle panel of Fig. S3. Thus, we create conditions that resemble the procedure postulated in the main text of the paper – but ‘à rebours’, i.e. effectively we end up with a removed spin site. This is because the spin exchange across the empty site $\tilde{J}$ is the same as the spin exchange $J$ elsewhere.
  4. 4.
    Finally, at time $\tau=\Gamma$ the RIXS x-ray is emitted and its momentum and energy measured, see right panel of Fig. S3. As the energy and momentum lost by RIXS is taken by the single spinon excited by RIXS, we can this way measure the single spinon dispersion.
Note that the above procedure does not depend on dimensionality. However, it requires finding a spin liquid based on a single-band Hubbard and that possesses a fine-tuned core-hole potential $U_{c}$ at the chosen indirect RIXS edge (although overall its value is reasonable Ament et al. (2011); Jia et al. (2016)). Finally, both pumps have to be perfectly aligned in time with the RIXS probe (incl. the core-hole lifetime).
We may wonder whether this way we can indeed excite just single spinons in higher dimensions – since removing a single site from e.g. a 2D square lattice also locally changes the topology of the lattice (which can be probed e.g. by traversing a loop around the created vacancy). While we leave this question to future studies, the point is that the presented scheme clearly excites single spinons in 1D and hence one should investigate in the future whether its natural generalisation to higher dimensions also predominantly excites single spinons – or there are also other excitations involved.
Last but not least, we note that instead of using RIXS, an alternative way to generate the coupling $\tilde{J}\sim J$ across the emptied site is to use the polarizability of ions, as suggested in Fig S4. For simplicity of drawing, the panels illustrate the idea for a 1D chain, but it can be generalized straightforwardly to higher dimensions.
In this procedure, we assume that the original spin lattice is mirrored by another similar lattice, populated by polarizable ions with full atomic shells and a polarizability $\alpha$ (top panel). When one of the spins is removed by the PES process (bottom panel), the resulting uncompensated charge at that site creates a local electric field that polarizes the nearby polarizable ions. This lowers the total energy by an amount of order $\sum_{\delta}\alpha{\vec{E}(\delta)}^{2}/2$, where $\vec{E}(\delta)$ is the electric field of the uncompensated charge at the polarizable ion located $\delta$ sites away Sawatzky et al. (2009). As such, this energy can be controlled by the appropriate choice of the polarizable ions and of their positioning with respect to the spin lattice. In particular, it can be set to the value needed to ensure that spin exchange $\tilde{J}$ across the empty site is equal to the exchange between nearest-neighbour spins.
Figure S3: Cartoon of the pump and probe experiment that can excite a single spinon by ‘kicking a spin and gluing the spins together’, i.e. following a procedure that qualitatively mimics the one outlined in the main text of the paper: Left panel – steps 1 and 2: photoemission spectroscopy (PES) creates a single hole that is followed by a single core-hole created by the resonant-inelastic x-ray scattering (RIXS) at the same site and at time $\tau\ll 1/t$; Middle panel – step 3: during time $0<\tau<\Gamma$ the RIXS core-hole potential lowers the value to such an extent that the value of the spin exchange $\tilde{J}$ across the empty site is equal to the exchange between nearest neighbor spins (see text for further details); Right panel – step 4: At time $\tau\sim\Gamma$ the RIXS photon is emitted, its energy and momentum is measured, and the core-hole filled. Note that in this simplified cartoon picture the pumps and the probe are aligned at the same lattice site $j$ – while in reality the two are aligned at the same momentum $k$ (see text for further details).
Figure S4: Cartoon showing how one can ‘kick a spin and glue the spins together’ using the photoemission and the polarisable ions: Spin lattice and the mirrored lattice that is populated by polarizable ions (top panel); When one of the spins is removed by the photomeission process, the resulting uncompensated charge at that site creates a local electric field that polarizes the nearby polarizable ions and lowers the total energy Sawatzky et al. (2009) (bottom panel). Note that the latter can be set to the value needed to ensure that spin exchange $\tilde{J}$ across the empty site is equal to the exchange between nearest-neighbour spins (see text for further details).